OpenStudy (anonymous):

Two complementary angles are in the ratio 3:6 find the measure of each angle.

hero (hero):

30 and 60

OpenStudy (anonymous):

how did you do that

hero (hero):

3/6 = 1/2 30/60 = 1/2 30 + 60 = 90

hero (hero):

Since 3/6 and 30/60 have the same ratio and since 30 + 60 = 90, we know that these are the correct angles

OpenStudy (anonymous):

ok thxs